a machinist was tasked to insert an aluminum piston into a steel cylinder. the inside diameters of the cylinder and of the piston are 0.500 m and 0. 4998 m, respectively, at 25 degree Celsius. To what temperature must the be heated to obtain a perfect fir?


1
Expert's answer
2022-01-27T13:17:44-0500

Gives

"L=L_0(1+\alpha(T-T_0))"


"500=0.4998(1+1.1\times(T-25))"

"\frac{500}{0.4998}-1=1.1\times(To-25)\\1000.40-1=1.1\times(T_0-25)"

"\frac{999.40}{1.1}=T_0-25"

"T_0-25=908.54=933.54\u00b0C"